OpenStudy (anonymous):

Given f(x)=3x^2+18x+6 where is its vertex?

Nnesha (nnesha):

formula to find x coordinate of vertex \[\huge\rm x=\frac{ -b }{ 2a }\] and then substitute x in f(x) function for x value to find y

Nnesha (nnesha):

a and b are \[\huge\rm Ax^2+Bx+C=0\] a=leading coefficient b=middle term

OpenStudy (anonymous):

so just solve it first

Nnesha (nnesha):

yes right first find x-coordinate

OpenStudy (anonymous):

21x?

Nnesha (nnesha):

how did you get that ?

Nnesha (nnesha):

a= ?? b= ?

OpenStudy (anonymous):

a=3x^2 b=18x

Nnesha (nnesha):

just the coefficient

Nnesha (nnesha):

so a=3 b=18 now use that formula that i posted bove

OpenStudy (anonymous):

-3

Nnesha (nnesha):

nvm yes

Nnesha (nnesha):

now to find y coordinate of vertex substitute all x in f(x) function for -3

OpenStudy (anonymous):

so its -3, -21

Nnesha (nnesha):

yes right